# AMEX: Android Multi-annotation Expo Dataset for Mobile GUI Agents

### v1

@@ 3.4 Level III: Instructions with GUI-Action Chains
Our focus for GUI agents is predominantly on third-party applications rather than system-built apps. We select an average of three apps per category from the Google Play Store. Due to the significant increase in app functionalities over the past years, we estimate that a sample size of approximately 50 apps can adequately represent most common use scenarios.
The instruction generation process comprises three phases. Initially, human annotators create 5-10 complex instructions for each target app, considering its specific purposes and capabilities. These initial instructions, combined with relevant app metadata collected online, serve as input for ChatGPT. It then generates a larger set of 80-100 instructions that exhibit similar structure and intent to the human-provided examples. For example, if a human annotator provides the instruction “Open Google Maps. Find the shortest route from the current location to Empire State Building by car drive,” ChatGPT might produce “Open Google Maps. Find the fastest route from the current location to Rockefeller Center by public transit.” However, due to ChatGPT’s limitations on understanding real-world constraints, human filtering is adopted to remove any unreasonable or impractical instructions. Notably, instructions within AMEX are more complex than those in the AitW dataset, with an average step count of 12.8, nearly double that of AitW.
@@
We define our action space for stepwise GUI operations similarly to AitW: {TAP, SCROLL, TYPE, PRESS_BACK, PRESS_HOME, PRESS_ENTER, TASK_COMPLETE, TASK_IMPOSSIBLE}. TAP actions are characterized by identical touch and lift coordinates, while SCROLL actions involve distinct touch and lift coordinates. TYPE actions are annotated with a type_text attribute specifying the input text. The three PRESS actions correspond to system-level button presses (back, home, enter). TASK_COMPLETE and TASK_IMPOSSIBLE serve as terminal flags for instructions. Additionally, for instructions that involve information query (e.g., “What is the lowest price of the men’s belt?”), we associate the TASK_COMPLETE action with a region of interest, which is defined as the bounding box of the area on the current screenshot, where the answer is expected to appear (see examples in Appendix A.3). This comprehensive action space allows us to fully simulate a wide range of typical use cases.
The collection tool is designed to simultaneously record the screen and action pair together with the XML file. The details of GUI-action chain collection can be found in Appendix A.1.4.
3.5 AMEX Statistics and Test Set
Table 2: AMEX statistics
# Screenshots # Interactive
Elements # Functionalities # Instructions # Avg. Steps # Apps
103,684 1,658,439 711,823 2,946 12.8 110
Figure 4: (a) The number of screenshots by the number of GUI clickable elements. (b) The number of instructions by the number of GUI operations steps. (c) The word cloud of the instructions.
Data statistics. Table 2 lists the statistics of the AMEX dataset. AMEX has approximately 104K screenshots, 1.6M GUI interactive elements for grounding, 712K GUI element functionality descriptions, and 3K instructions. Figure 4 illustrates the distribution of GUI interactive elements per screen and GUI operation steps along with the word cloud of the instructions.
Test set. Instructions and their annotated screen-action chains from 9 out of 55 apps are chosen as the test set, including those on Booking, Gmail, NBC News (News), SHEIN, Citymapper (CM), Microsoft To Do (ToDo), Signal, Yelp and YouTube Music (Music), covering different scenarios. The number of test instructions is 362, around 10% of the total instructions.
The evaluation of AMEX test set is similar to the existing dataset AITW [19]. We perform the action matching to get the partial action match score. The partial action matching score is the number of correct actions divided by the episode length [14]. To ensure the correctness of the test set, we carefully review the trajectory annotations in the unseen test set, where we drop the former step of PRESS_BACK, which indicates the wrong operation of the annotator. However, we keep the PRESS_BACK step to evaluate the agent’s self-correction ability.
4 SPHINX-GUI Agent
Integrating data from AMEX with the existing AitW dataset, we train the SPHINX-GUI Agent tailored for GUI-related tasks as the baseline for AMEX dataset. This Agent is initialized from SPHINX [7] and is specifically trained to generate actions to manipulate the GUI effectively, which are different from the more flexible output formats typical of standard MLLM tasks such as QA. To incorporate GUI-specific intelligence into SPHINX, we have implemented several key modifications:
Instruction Dataset Construction. We transform the AMEX data and the publicly available AitW dataset into a VQA format conducive to instruction understanding, thereby adapting the MLLM as a GUI agent. To embed GUI-specific knowledge within the MLLM, we developed four distinct VQA tasks, detailed in the instruction templates provided in Table 3.
Table 3: Task template examples.
Tasks Examples of Task Templates Source
Screen Description
User: Provide a one-sentence caption for the provided GUI image.
Agent: This page is a film section of the IMDb app, displaying a list of movies or shows with sorting and filtering options.
AMEX
Element Grounding
User: Identify all clickable elements and provide their 2D bounding boxes.
Agent: List of Elements with BBoxes.
AMEX
Functionality Description
User: What is the function of the icon at the coordinates: BBox.
Agent: Click to expand or collapse the sort options.
AMEX
GUI-Action Prediction
User: Please execute the task described by the following instruction: Delete the first email in the inbox.
Agent: CLICK[[0.59, 0.92]].
AMEX & AitW
Structured Representation. To effectively guide the agent in task completion, we incorporate the goal into the input prompt. To prevent redundant actions, we also include a history of actions formatted as $X_{\text{history}}=\left[a_{t-k},\ldots,a_{t}\right]$, where each $a$ denotes an action type, accompanied by a touch point for actions such as SCROLL or TAP. The parameter $k$ denotes the length of the history. Action trajectories from both AitW and AMEX datasets are converted into this uniform representation.
Given the extensive nature of GUI screenshots and the small relative size of some target elements within the screen, we adopt the “any resolution” approach from SPHINX [7]. This method initially partitions the input image into sub-images, then encodes separately. The LLM processes all the resulting visual tokens to ensure comprehensive understanding and interaction capabilities. We follow all other default training receipts in [7] to obtain the final agent.
@@ 5 Experiments
5.1 Evaluation Setup and Compared LVLMs
We evaluate different GUI agents including SPHINX Agent developed in Section 4 on AitW and our AMEX dataset respectively. For AitW, we simultaneously train SPHINX Agent on all the subsets and then assess all test sets. We downsample GoogleApps subset to $10\%$ to avoid data imbalance. The training and test split setting follows the one in [19]. For AMEX, we evaluated our SPHINX agents on the test set as described in Section 3.5. Our models are benchmarked against several agents including Auto-UI [30], SeeClick [3], CogAgent [10] and CoCo-Agent [17].
5.2 Experiment Results
The experiment outcomes for AitW and AMEX are presented in Tables 4 and 5 respectively. As indicated in Table 4, SPHINX Agent (SphAgent) trained on AitW exhibit competitive performance on the AitW dataset relative to the baselines. Notably, the introduction of the AMEX dataset enhanced the overall performance by approximately 2.5%, with significant gains observed in the “General” (5%), “Single” (5%) and “WebShopping” (2%) tasks, which indicates the strong complement from AMEX dataset. The GUI element functionality descriptions are served as the “Single” step instructions which strongly promote the performance in the “Single” tasks, while 3K instructions from general third-party apps also boost the performance in the “General” tasks. “WebShopping” is the category with the most complicated tasks in AitW, which also benefits from the complex instructions provided by AMEX.
Table 4: Experiment results on AitW. + indicates CoCo-Agent takes screen element OCR layouts as additional inputs.
@@ 5.2 Experiment Results
Agent # Params Training Data General Install G-Apps Single WebShopping Overall
Auto-UI 5B AitW 68.24 76.89 71.37 84.58 70.26 74.27
[…]
CoCo-Agent+ 7B AitW 70.96 81.46 76.45 91.41 75.00 79.05
SphAgent 7B AitW 68.2 80.5 73.3 85.4 74 76.28
SphAgent 7B AitW + AMEX 73.1 80.5 73.4 90.8 75.8 78.72
@@
Table 5: Experiment results on AMEX
@@
Agent Training Data Gmail Booking Music SHEIN News CM ToDo Signal Yelp Overall
[…]
SphAgent AitW 32.1 45.9 46.1 35.1 48.3 61.1 55.9 43.3 42.9 45.63
SphAgent AMEX 61.7 68.2 77.7 72.0 71.9 64.6 79.6 71.3 69.6 70.71
SphAgent AitW + AMEX 62.4 68.1 76.3 71.9 68.6 67.3 77.6 66.0 64.1 69.14
@@ 5.3 Cross-Domain Experiments
We also examine the generalization capabilities of our agent across different domains. Specifically, the agent trained exclusively on the AitW dataset is tested on the AMEX dataset. The results, as detailed in the first row of Table 5, reveal significant findings. Agent trained only on AitW, shows satisfactory performance on the AitW dataset itself. However, when evaluated on the AMEX dataset, there is a notable decrease in performance. This performance degradation can likely be attributed to several factors: 1) Higher complexity in AMEX. The AMEX dataset generally involves longer operational sequences compared to those in AitW, presenting a higher complexity level that the agent might not have been exposed to during training. 2) Domain gap. There are substantial differences in the visual textures and language instructions between the two datasets. These variations could hinder the agent’s ability to effectively transfer its learned knowledge to a new, unfamiliar domain. Additionally, we observed a significant performance degradation for the SeeClick agent, originally trained on AitW. Beyond the previously mentioned reasons, this decline may also stem from the fact that the original SeeClick model was designed to process only $224\times 224$ image inputs. In contrast, the AMEX dataset comprises high-resolution images and requires interactions with tiny icons. The high-definition images in AMEX introduce new challenges for GUI agents.
5.4 AitW Human Evaluation Results
MMR SphAgent 90.91 83.65 87.50 87.68 82.39 86.43
AitW Anno 88.18 75.00 82.69 84.78 75.57 81.25
Random SphAgent 92.31 93.75 94.33 94.83 94.55 93.95
AitW Anno 94.23 90.34 93.62 93.97 95.00 93.43
As specified in AitZ [32], several types of error cases are identified in the AitW test set (see Figure5). To assess the unreliability of the original AitW test set and its evaluation methods, we propose AitW-HE (Human Evaluation), a refined subset of the AitW test set. Human annotators evaluate two subsets derived from the original test set. One subset is randomly chosen from episodes where SphAgent receives low scores (Mis-Match Random), i.e., about 66 score compared to the overall 78.72 shown in Table 4, indicating a high mismatch between inference results and AitW annotations. The other subset is randomly selected from the remaining episodes (Random). Annotators first filter out repeated and redundant screenshots (see Figure 5(a)). For each remaining screenshot, they then record whether the SphAgent-inferred action and the original AitW annotation are correct. Figure 6 illustrates cases where annotators mark both the inferred action and the original annotation as correct, even though they interact with different elements.
(a) Repeating useless screenshots and actions.
(b) Wrong task complete.
Table 6 presents the accuracy scores from human evaluation. The table indicates that the low SphAgent scores in the MMR subset are primarily due to unsatisfactory annotations (AitW Anno), which are used as ground truth during the evaluation. Both SphAgent and AitW Anno in the “MMR” subset have lower scores than those in the “Random” subset, highlighting that the tasks in the MMR subset are relatively more challenging. Furthermore, SphAgent achieves better human evaluation results than the original annotations, demonstrating its effectiveness and close alignment with human judgment. Comparing the results from Table 4 and Table 6, the notable differences in overall scores (i.e., 78.72%, 86.43%, and 93.95%) underscore the unreliability and misleading nature of the AitW test set and its evaluation methods.
5.5 Ablation Study on Multi-level Annotations
To validate the effectiveness of the multi-level annotations proposed in the AMEX dataset, we conducted an ablation study. As shown in Table 7, the agent achieved a performance gain of approximately 1.0 when aided by L1 annotations. When trained with L1 and L3 annotations, the agent exhibited a performance gain of about 1.6 compared to the baseline. The gain from L2 annotations is approximately 2.3. Furthermore, training the agent with the complete set of annotations resulted in a performance gain of approximately 3.5. The ablation study results demonstrate that each level of annotations in AMEX enhances the final agent’s performance.
Table 7: Ablation study of each level of annotations.
Tasks Gmail Booking Music SHEIN NBC CM ToDo Signal Yelp Avg
L1 L2 L3
45.9 64.5 74.4 71.8 70.3 67.4 79.3 64.9 66.3 67.2
58.6 68.3 74.4 68.8 68.1 61.1 79.3 73.2 67.4 68.8
60.3 66.1 77.2 70.2 72.0 64.0 78.1 72.3 65.7 69.5
61.7 68.2 77.7 72.0 71.9 64.6 79.6 71.3 69.6 70.7
@@ 6.1 Limitations and Future Work
Multi-lingual    Most existing datasets are limited to English, with UGIF [23] being a notable exception, as it includes instructions and screenshots in eight languages. The AMEX dataset contains a small number of screenshots in Chinese and Spanish, primarily due to strict registration and login requirements for Chinese apps and a lack of expertise in other languages. Future work should incorporate multi-lingual screenshots, functionalities, and instructions to create a more robust and comprehensive multi-lingual environment for GUI agents.
Evaluation    The current evaluation method employed in both AMEX and AitW is straightforward: the model predicts the current action based on given instructions, previous operations, and screenshots. However, this evaluation approach has significant limitations in real-world scenarios. It fails to account for loading pages, waiting times for page loading, etc. An ideal evaluation would be an “Arena” (a mobile version of the environment in WebArena [35]) — an independent local virtual device equipped with Android and static background apps. This setup would ensure consistent results for the same queries, avoiding variations due to differing dates, prices, and other dynamic data.
@@ A.1 Pipeline Details
A.1.1 Autonomous script details
@@
A.1.2 GUI clickable element filtering
@@
A.1.3 GUI screen and element functionality description collection details
We apply the Set-of-Mark (SoM) [27] technique when using GPT-4o as the description generator. The SoM technique is a visual prompting method designed to enhance the visual grounding capabilities of large multimodal models (LMMs), such as GPT-4V, by overlaying visual marks on image regions. This involves partitioning an image into semantically meaningful regions and adding distinct marks (e.g., alphanumeric characters, masks, or boxes) to these regions. It demonstrates significant improvements in precision and accuracy over traditional prompting methods and other state-of-the-art models. Figure 8 shows the screenshot with SoM technique.

### v2

@@ 3.4 Level III: Instructions with GUI-Action Chains
Table 1 illustrates that most instruction-based datasets provide an average number of operation steps below 7, which implies most of the instructions are simpler than tasks in real-world. For example, users often apply filters or sort by different attributes when searching, which causes more complex instructions and longer action chains. In order to address this situation, we collect a set of complex instructions with GUI-action chains on popular apps from Google play store.
@@
We define our action space for stepwise GUI operations similarly to AitW: {TAP, SCROLL, TYPE, PRESS_BACK, PRESS_HOME, PRESS_ENTER, TASK_COMPLETE, TASK_IMPOSSIBLE}. TAP actions are characterized by identical touch and lift coordinates, while SCROLL actions involve distinct touch and lift coordinates. TYPE actions are annotated with a type_text attribute specifying the input text. The three PRESS actions correspond to system-level button presses (back, home, enter). TASK_COMPLETE and TASK_IMPOSSIBLE serve as terminal flags for instructions. Additionally, for instructions that involve information query (e.g., “What is the lowest price of the men’s belt?”), we associate the TASK_COMPLETE action with a region of interest, which is defined as the bounding box of the area on the current screenshot, where the answer is expected to appear (see examples in Appendix A.3). This comprehensive action space allows users to fully simulate a wide range of typical use cases. Detailed collection process and instruction generation are in Appendix A.1.5.
@@ 4 Experiments
AMEX serves as a supplementary dataset for other large-scale instruction-based datasets. To test the effectiveness of AMEX, we directly finetune SPHINX models Liu et al. (2024) without any tricks and conduct evaluation on three other benchmark datasets. We choose ScreenSpot Cheng et al. (2024) as the benchmark of our agent’s element grounding ability and AitW and AndroidControl as the benchmarks of our agent’s GUI-control ability.
4.1 SPHINX Agent
Our SPHINX agent, SphAgent for short, is initialized from SPHINX Liu et al. (2024) and it is finetuned for the GUI-control tasks and element grounding tasks. For both tasks, we train the model in a pure vision-based principle, which means the model understands the environment by only screenshots, without any other supplementary information such as accessibility trees or view hierarchy, due to the fact that many apps don’t support or only provide very little accessibility tree or view hierarchy element information. We consider vision-based agents are more applicable in more various use cases.
Table 3: ScreenSpot mobile subset experiment results.
Model Size Icon / Widget
Fuyu 8B 1.3%
CogAgent 18B 24.0%
SeeClick 9.6B 52.0%
Qwen2-VL 7B 60.7%
GPT-4V w. OmniParser - 57%
SphAgent 7B 72.6%
4.2 ScreenSpot
Table 4: Experiment results on AndroidControl. “Click & Long Press” is specially computed to show the improvements on these actions. The best results are in bold for different subsets and levels.
Agent # data Training Data Task
Level IDD Category
Unseen App
Unseen
Task
Unseen Overall Click &
Long Press
SphAgent 178K AndroidCtrl High Lv. 58.3 40.6 39.4 52.2 49.8 34.1
Low Lv. 79.8 71.0 71.1 83.5 75.8 53.0
SphAgent 178K 70% AndrCtrl High Lv. 60.4 41.7 43.4 59.3 52.8 37.3
+ 10% AMEX Low Lv. 81.9 73.2 72.6 87.5 77.7 56.7
SphAgent 712K AndroidCtrl High Lv. 70.5 51.6 55.6 70.2 61.7 50.8
+ AMEX Low Lv. 88.9 81.8 76.9 92.6 84.2 67.0
ScreenSpot Cheng et al. (2024) is a benchmark dataset which contains over 600 UI screenshots from mobile devices, desktops and websites. It is designed to evaluate the model’s element grounding capability, where each human-annotated functionality corresponds to an interactive element. We train SPHINX agent on AMEX level 1 and level 2 data to evaluate the element grounding capability.
Table 3 lists the evaluation results on the ScreenSpot mobile subset. The models Bavishi et al. (2023); Hong et al. (2024); Cheng et al. (2024); Wang et al. (2024) in first four rows are LVLMs specially trained on GUI data. OmniParser Wan et al. (2024) is a combination of an icon detection model, an icon description model and an OCR module, which serves as a screenshot parser for other LVLMs such as GPT-4V. During evaluation, our SphAgent surpasses four LVLMs by a large margin on the “Icon / Widget” subset without any training tricks, proving that the functionality understanding can largely boost the performance of GUI grounding for agents. The above results and comparisons strongly prove the effectiveness and applicability of AMEX in GUI element grounding tasks.
4.3 AndroidControl
AndroidControl Li et al. (2024) is a large-scale benchmark dataset to evaluate the agent’s GUI-control ability to complete tasks on general apps (see more detailed description of the dataset and splits in Appendix A.2.1). We evaluate three SphAgents trained on (i) only AndroidControl, (ii) random 70% AndroidControl data mixed with random 10% AMEX level 1 and 2, (iii) a mix of AndroidControl and AMEX level 1 and 2. The (ii) experiment ensures the same number of data point as in (i) experiment. The action space of AndroidControl is different from AMEX level 3 and AitW, so we only use level 1 and level 2 data during the experiments (ii) and (iii).
Table 4 lists the evaluation results on AndroidControl test set. The evaluation results of experiment (i) and (iii) strongly support the effectiveness of AMEX level 1 and level 2. Adding full data of AMEX level 1 and level 2 leads to an average 10% overall performance gain and strongly improves the “click” and “long press” actions by more than 14%. Also, the comparative results from experiment (i) and (ii) shows that even using the same number of data points, replacing down-stream instructional task data with environment understanding data would also lift the performance of agents on both low-level and high-level at an average of 2.5%. This performance gain indicates the effectiveness of our multi-level annotations.
4.4 AitW
AitW is another large-scale benchmark dataset to evaluate the agent’s GUI-control ability to complete task goals. We evaluate three SphAgents trained on (i) only AitW, (ii) a mix of AitW and AMEX. Since the GUI-action chains in AMEX share the same action space with AitW, we utilize all three levels of AMEX in the mixed data.
Table 5 lists the evaluation results on AitW test set, which has five subsets for different types of tasks. SphAgent trained on the mixed data (AitW + AMEX) achieves the highest overall accuracy among all agents, and when compared to the SphAgent trained on only AitW, the accuracies in “General” and “Single” categories have a notable gain of around 5% and the overall accuracy lifts by 2.5%. That the gains in other categories are not remarkable is possibly due to the mis-aligned element-wise annotation and unregulated instruction operation annotation. In Appendix A.2.2 we further explore the dataset evaluation and explain the reason why the evaluation results might be misleading.
Table 5: Experiment results on AitW.
@@ 4.4 AitW
Agent Training Data General Install G-Apps Single WebShopping Overall
SphAgent AitW 68.2 80.5 73.3 85.4 74 76.28
SphAgent AitW + AMEX 73.1 80.6 73.4 90.8 75.8 78.72
SphAgent AMEX 51.5 55.6 57.1 61.9 55.1 56.2
@@
Table 6: Experiment results on AMEX. The test apps are excluded during the training of SphAgent on AMEX.
@@
Agent Training Data Gmail Booking YT-Music SHEIN NBC CM ToDo Signal Yelp Overall
SphAgent AitW 32.1 45.9 46.1 35.1 48.3 61.1 55.9 43.3 42.9 45.6
SphAgent AMEX 63.7 67.4 76.1 70.0 68.5 62.7 78.6 68.2 68.9 69.3
@@ 4.5 Cross-Domain Experiments
Since AMEX Level 3 shares the same action space only with AitW, we conducted two experiments: (i) training SphAgent on AMEX and testing it on AitW, and (ii) training SphAgent on AitW and testing it on AMEX. The results, as shown in the last row of Table 5, indicate poor performance when the agent is trained only on AMEX and tested on AitW. Similarly, Table 6 demonstrates poor performance when the agent is trained on AitW and tested on AMEX. This significant discrepancy in evaluation results is likely due to the domain gap between the two datasets. First, the styles of collecting GUI-action chains differ between the two datasets. For instance, annotators in AitW tend to directly search for apps, whereas annotators in AMEX prefer navigating through the app library to locate apps. This inconsistency in interaction strategies poses a challenge for the agent, which evaluates tasks based on static frames rather than dynamically adapting to different usage patterns. Second, although both datasets share the same action space, they differ in terms of visual appearance and versioning. The apps in AMEX are updated to their latest versions, while those in AitW are based on older versions. These versioning differences lead to changes in the visual appearance of GUI elements (e.g., color schemes, icons, or button designs), which can negatively impact the agent’s ability to recognize and interact with them.
@@ 5.1 Limitations and Future Work
Multi-lingual
Most existing datasets are limited to English, with UGIF Gubbi Venkatesh et al. (2024) being a notable exception, as it includes instructions and screenshots in eight languages. The AMEX dataset contains a small number of screenshots in Chinese and Spanish, primarily due to strict registration and login requirements for Chinese apps and a lack of expertise in other languages. Future work should incorporate multi-lingual screenshots, functionalities, and instructions to create a more robust and comprehensive multi-lingual environment for GUI agents.
Online Evaluation
Due to the limitation that SPHINX is not deployable, we encountered issues trying online evaluation such as AndroidWorld Rawles et al. (2024a). In future work, online dynamic evaluation is more accurate and simulate the real-world agent usage.
Multi-Platform
Android is an open-source OS and the community provides well-developed tools for interactions and emulations. iOS is another widely used mobile device OS but it is restricted and hard to run test on. In the future, a dataset extended to other OS or platforms would provide more general information.
@@ A.1 Pipeline Details
The raw data is collected using Android emulators, specifically Android Virtual Device and Genymotion [^1] https://www.genymotion.com/. We developed tools to autonomously perform emulator operations and record human actions. These tools leverage Appium [^2] https://appium.io/docs/en/latest/, an open-source, cross-platform test automation tool for Android, iOS, and web apps, under Apache license. Appium collects device screenshots during operations, each of which also corresponds to an XML file recording the basic screen layout with element attributes, such as bounding boxes and in-app descriptions.
A.1.1 Collection details
As mentioned in Section 3.1, the collection of screenshots and their XML data utilizes two methods. Here, we provide a more detailed pipeline for executing an autonomous script to perform operations on an Android emulator. The script is designed to traverse an app in an unconstrained manner and collect data. It performs actions (see Appendix A.1.2) at regular time intervals to allow each page to fully load. Then the script captures the current screenshot along with the corresponding XML file. Upon the raw data (yellow part in Figure 2), we extract bounding boxes for interactive elements from the XML associated with each screenshot. Then annotators manually inspect each screenshot, identifying and selecting only the elements that are actually visible within the interface (see Appendix A.1.3 for detailed reasons).
A.1.2 Autonomous script details
@@
A.1.3 GUI clickable element filtering
@@
A.1.4 GUI screen and element functionality description collection details
We adopt GPT-4o and Gemini 1.5 Pro to describe screens and interpret element functionalities separately. Given the compact and dense nature of GUI elements, we apply Set of Mark (SoM) techniques Yang et al. (2023) to boost GPT and Gemini’s capability for visual localization. The coordinates of these elements are obtained and cleaned in the preceding phases (Section 3.2). Additionally, recognizing that some elements possess abstract functionalities that are challenging to discern, we supplement the prompts for GPT and Gemini with in-app descriptions extracted from XMLs to enhance their comprehension of the screen. After the generation, we further utilize GPT-4o to compare the functionalities of the same element generated by two models and keep them if two functionalities have the same meaning. Finally, human annotators verify the quality and the functionality accuracy rate is above 97%.
We apply the Set-of-Mark (SoM) technique when using GPT and Gemini as the description generator. The SoM technique is a visual prompting method designed to enhance the visual grounding capabilities of large multimodal models (LMMs), such as GPT-4V, by overlaying visual marks on image regions. This involves partitioning an image into semantically meaningful regions and adding distinct marks (e.g., alphanumeric characters, masks, or boxes) to these regions. It demonstrates significant improvements in precision and accuracy over traditional prompting methods and other state-of-the-art models. Figure 5 shows the screenshot with SoM technique.
The most recurring mistake is the mis-interpretation of the content description. For instance, some “compound” elements, which may contain several text, rating stars, image, etc., have couple of content descriptions. Those mixed content descriptions may cause the misunderstanding of the real functionality of the compound element. In IMDB app, a row of movie entry sometimes would consist of a poster, a text name, ratings of the movie, and the price of renting. The desired functionality of the entry could be "view the detail page of the movie xxx", while GPT might generate "view the price for renting the movie xxx", which is not exactly what we want. However, we use two LLMs to double check their generations on the same element functionality and we only keep the elements with the same functionalities generated by two LLMs.